OpenStudy (precal):

Find y' y=(xsinx)/(1+cosx)

OpenStudy (precal):

I think I do quotient rule

OpenStudy (anonymous):

yes i think you need that one

OpenStudy (anonymous):

also produce rule for the numerator, but when you are done you should just get \[y'=\frac{x+\sin(x)}{\cos(x)+1}\]

OpenStudy (precal):

Thanks